# The early r-process nucleosynthesis scenarios

### v1

@@ I Introduction
The rapid neutron-capture process (r-process), in which intermediate-mass elements capture tens to about 200 neutrons, occurs in dynamically evolving, extremely neutron-rich sites, leading to the nucleosynthesis of the heaviest elements (e.g., Thielemann & Cowan 2026, for a recent review). There are several r-process scenarios: (1) the merger of two neutron stars, the NS-NS merger scenario (e.g., Goriely et al. 2011; Wanajo et al. 2014; Beniamini et al. 2016a, b; Beniamini & Piran 2024; Ji et al. 2016; Metzger 2017; Banerjee et al. 2020; Dvorkin et al. 2021; van de Voort et al. 2022; Maoz & Nakar 2025; Qiumu et al. 2025; Rastinejad et al. 2025); (2) the magnetorotational supernova scenario, i.e., core-collapse supernovae (CCSNe) that are powered by a fixed-axis pair of jets (e.g., Winteler et al. 2012; Nishimura et al. 2015, 2017; Halevi & Mösta 2018; Reichert et al. 2021, 2023; Yong et al. 2021; Liu et al. 2025); (3) the NS winds scenario, i.e., a wind from the newly born NS in a CCSN (e.g., Prasanna et al. 2024), possibly limited to the first r-process peak, i.e., the weak r-process (e.g., Wang & Burrows 2023, 2024), but with the third peak in the presence of magnetars (Prasanna et al., 2025); (4) the common envelope jets supernova (CEJSN) r-process scenario (e.g., Papish et al. 2015; Grichener & Soker 2019a; Grichener et al. 2022; Grichener & Soker 2022; Grichener 2023, 2025; Soker 2025); (5) collapsars, i.e., core-collapse supernova that forms a black hole at the center (e.g., Siegel et al. 2019, 2022; Brauer et al. 2021; Issa et al. 2025; Leicester et al. 2025; Mumpower et al. 2025), which probably requires significant magnetic fields (Just et al. 2022); (6) magnetar giant flares (e.g., Cehula et al. 2024; Negro et al. 2025; Patel et al. 2025a, b); (7) accretion induced collapse (AIC) of an ONeMg white dwarf (e.g., Batziou et al. 2025; Cheong et al. 2025; Combi et al. 2025; in this study, I include the merger of two white dwarfs, i.e., merger induced collapse, under AIC).
@@
As of today, observations directly and clearly support the NS-NS merger scenario, indicating the presence of very heavy radioactive elements (e.g., Chornock et al. 2017; Kasen et al. 2017; Pian et al. 2017; Watson et al. 2019; Kasliwal et al. 2022; Watson et al. 2019; Domoto et al. 2022; Levan et al. 2024), and possibly in one magnetar giant flare (Patel et al., 2025a). However, several papers suggest that the NS-NS merger scenario may not be the sole r-process site (e.g., Waxman et al. 2018; Côté et al. 2019; Ji et al. 2019; Kobayashi et al. 2020, 2023; Holmbeck & Andrews 2024; Thielemann & Cowan 2026 for a review). Particularly, studies claimed that the lanthanide fraction (relative to total r-process yield) of kilonova GW170817/AT 2017gfo is much below the observed value, and can be more than an order of magnitude below the observed values: (Waxman et al., 2018), (Ji et al., 2019), and (Gillanders et al., 2025); most metal-poor stars have , as I present in Figure 1 adapted from Jin & Soker (2024). Many papers discuss the need for two or more r-process sites to explain the Milky Way r-process abundance (e.g., Wehmeyer et al. 2015; Grichener & Soker 2019b; Haynes & Kobayashi 2019; Molero et al. 2021; Tsujimoto 2021; Farouqi et al. 2022; Naidu et al. 2022; Yamazaki et al. 2022; Storm et al. 2025). Rastinejad et al. (2025) who study the NS-NS merger scenario, find that there is room for more r-process sites in addition to the NS-NS merger.
@@
The r-process scenarios are a hot topic, with numerous recent papers addressing this topic (e.g., Hayes et al. 2023; Placco et al. 2023; Anoardo et al. 2025; Bishop et al. 2025; Hansen et al. 2025; Hirai et al. 2025; Kobayashi 2025; Maoz & Nakar 2025; Molero et al. 2025; Palla et al. 2025; Santarelli et al. 2025; Storm et al. 2025; Tian et al. 2025; Xie et al. 2024, 2025). The goal of this paper is to examine actively studied r-process sites in light of recent studies of r-process nucleosynthesis in metal-poor stars, particularly in the early Universe. Namely, I address one question: Which r-process scenarios can contribute to the r-process in extremely metal-poor stars? Although this study presents no new calculations, it is the only study to systematically compare all actively studied r-process scenarios, omitting none; this allows me to conclude which scenarios are likely to contribute to r-processes in the early Universe. Therefore, it offers a timely summary of new results from 2025, and a valuable comparison of theoretical scenarios with observed r-process element properties in the early Universe. The conclusions, summarized in Section IV, have significant implications for future theoretical studies of r-process scenarios.
@@ II.1 Large abundance scatter
From very early times of having [Fe/H], the abundance of r-process elements, like Eu, shows very large scatter, as I reproduce in Figure 2 based on Kobayashi et al. (2020). This shows that there is no averaging over many r-process nucleosynthesis events at those early times. Therefore, each event must produce a relatively large mass of r-process elements. From [Fe/H], the scatter decreases with increasing metallicity (e.g., Ou et al. 2025), indicating the averaging over more and more r-process events; above , iron production in type Ia supernovae also reduces the scatter. Henderson et al. (2025) studied 89 stars in the globular cluster M15 and found high Ba, La, and Eu dispersions in the first generation of stars. They conclude, under the assumption that the r-process events that caused the abundance dispersions were born with the first population of stars in M15, that the r-process site must have a short delay time.
@@ III Scenarios comparison
In Table 1, I compare the seven scenarios listed in Section I with respect to their ability to explain the observations related to the early r-process in the Universe described in Section II. I consider two types of r-process sites (Section II.2): one that produces the light elements in the early universe, with some correlation with iron production (I term it “Eu”+Fe), and the other that forms the third r-process peak. I also require that the scenarios operate with a short delay after the first CCSNe and account for the large scatter in r-process elements (Section II.1). A ‘’ symbol indicates that the scenario is not a main contributor to the nucleosynthesis site in the early universe. It does not imply that it does not occur. I cannot rule out that it occurs, but only that it is not the main contributer. Three of the scenarios predict that the r-process event occurs long after the CCSN; hence, in these scenarios, r-process elements are generally not correlated with iron production in the early Universe (before type Ia supernovae occur). These are indicated by ‘No’ in the third column. The other four occur alongside the CCSN explosion (or starting within seconds after the explosion).
@@
| AIC | [Co25] | No | () Too long delay ; () No Fe expected | () Too long delay |

### v2

@@ I Introduction
The rapid neutron-capture process (r-process), in which intermediate-mass elements capture tens to about 200 neutrons, occurs in dynamically evolving, extremely neutron-rich sites, leading to the nucleosynthesis of the heaviest elements (e.g., Thielemann and Cowan 2026, for a recent review). There are several r-process scenarios: (1) the merger of two neutron stars, the NS-NS merger scenario (e.g., Goriely et al. 2011; Wanajo et al. 2014; Beniamini et al. 2016a, b; Beniamini and Piran 2024; Ji et al. 2016; Metzger 2017; Banerjee et al. 2020; Dvorkin et al. 2021; van de Voort et al. 2022; Maoz and Nakar 2025; Qiumu et al. 2025; Rastinejad et al. 2025); (2) the magnetorotational supernova scenario, i.e., core-collapse supernovae (CCSNe) that are powered by a fixed-axis pair of jets (e.g., Winteler et al. 2012; Nishimura et al. 2015, 2017; Halevi and Mösta 2018; Reichert et al. 2021, 2023; Yong et al. 2021; Liu et al. 2025); (3) the NS winds scenario, i.e., a wind from the newly born NS in a CCSN (e.g., Prasanna et al. 2024), possibly limited to the first r-process peak, i.e., the weak r-process (e.g., Wang and Burrows 2023, 2024), but with the third peak in the presence of magnetars (Prasanna et al., 2025); (4) the common envelope jets supernova (CEJSN) r-process scenario (e.g., Papish et al. 2015; Grichener and Soker 2019a; Grichener et al. 2022; Grichener and Soker 2022; Grichener 2023, 2025; Soker 2025); (5) collapsars, i.e., core-collapse supernova that forms a black hole at the center (e.g., Siegel et al. 2019, 2022; Brauer et al. 2021; Issa et al. 2025; Leicester et al. 2025; Mumpower et al. 2025), which probably requires significant magnetic fields (Just et al. 2022); (6) magnetar giant flares (e.g., Cehula et al. 2024; Negro et al. 2025; Patel et al. 2025a, b); (7) accretion induced collapse (AIC) of an ONeMg white dwarf (e.g., Batziou et al. 2025; Cheong et al. 2025; Combi et al. 2025; Pitik et al. 2026b, a; in this study, I include the merger of two white dwarfs, i.e., merger induced collapse, under AIC).
@@
As of today, observations directly and clearly support the NS-NS merger scenario, indicating the presence of very heavy radioactive elements (e.g., Chornock et al. 2017; Kasen et al. 2017; Pian et al. 2017; Watson et al. 2019; Kasliwal et al. 2022; Watson et al. 2019; Domoto et al. 2022; Levan et al. 2024), and possibly in one magnetar giant flare (Patel et al., 2025a). However, several papers suggest that the NS-NS merger scenario may not be the sole r-process site (e.g., Waxman et al. 2018; Côté et al. 2019; Ji et al. 2019; Kobayashi et al. 2020, 2023; Holmbeck and Andrews 2024; Chen et al. 2025; Henderson et al. 2025b; Saleem et al. 2025; Thielemann and Cowan 2026 for a review). Particularly, studies claimed that the lanthanide fraction (relative to total r-process yield) of kilonova GW170817/AT (SN 2017gfo) is much below the observed value, and can be more than an order of magnitude below the observed values: (Waxman et al., 2018), (Ji et al., 2019), and (Gillanders et al., 2025); most metal-poor stars have , as I present in Figure 1 adapted from Jin and Soker (2024). Many papers discuss the need for two or more r-process sites to explain the Milky Way r-process abundance (e.g., Wehmeyer et al. 2015; Grichener and Soker 2019b; Haynes and Kobayashi 2019; Molero et al. 2021; Tsujimoto 2021; Farouqi et al. 2022; Naidu et al. 2022; Yamazaki et al. 2022; Storm et al. 2025). Rastinejad et al. (2025) who study the NS-NS merger scenario, find that there is room for more r-process sites in addition to the NS-NS merger.
@@
The r-process scenarios are a hot topic, with numerous recent papers addressing this topic (e.g., Hayes et al. 2023; Placco et al. 2023; Anoardo et al. 2025; Bishop et al. 2025; Hansen et al. 2025; Hirai et al. 2025; Kobayashi 2025; Maoz and Nakar 2025; Molero et al. 2025; Palla et al. 2025; Santarelli et al. 2025; Storm et al. 2025; Tian et al. 2025; Xie et al. 2024, 2025). The goal of this paper is to examine actively studied r-process sites in light of recent studies of r-process nucleosynthesis in metal-poor stars, particularly in the early Universe. Namely, I address one question: Which r-process scenarios can contribute to the r-process in extremely metal-poor stars? Although this study presents no new calculations, it is the only study to systematically compare all actively studied r-process scenarios, omitting none; this allows me to conclude which scenarios are likely to contribute to r-processes in the early Universe. Therefore, it offers a timely summary of new results from 2025, and a valuable comparison of theoretical scenarios with observed r-process element properties in the early Universe. The conclusions, summarized in Section IV, have significant implications for future theoretical studies of r-process scenarios.
@@ II.1 Large abundance scatter
From very early times of having [Fe/H], the abundance of r-process elements, like Eu, shows very large scatter, as I reproduce in Figure 2 based on Kobayashi et al. (2020). This shows that there is no averaging over many r-process nucleosynthesis events at those early times. Therefore, each event must produce a relatively large mass of r-process elements. From [Fe/H], the scatter decreases with increasing metallicity (e.g., Ou et al. 2025), indicating the averaging over more and more r-process events; above , iron production in type Ia supernovae also reduces the scatter. Henderson et al. (2025a) studied 89 stars in the globular cluster M15 and found high Ba, La, and Eu dispersions in the first generation of stars. They conclude, under the assumption that the r-process events that caused the abundance dispersions were born with the first population of stars in M15, that the r-process site must have a short delay time.
@@ III Scenarios comparison
In Table 1, I compare the seven scenarios listed in Section I with respect to their ability to explain the observations related to the early r-process in the Universe described in Section II. I consider two types of r-process sites (Section II.2): one that produces the light elements in the early universe, with some correlation with iron production (I term it “Eu”+Fe), and the other that forms the third r-process peak. I also require that the scenarios operate with a short delay after the first CCSNe and account for the large scatter in r-process elements (Section II.1). A ‘’ symbol indicates that the scenario is not a main contributor to the nucleosynthesis site in the early universe. It does not imply that it does not occur. I cannot rule out that it occurs, but only that it is not the main contributor. Three of the scenarios predict that the r-process event occurs long after the CCSN; hence, in these scenarios, r-process elements are generally not correlated with iron production in the early Universe (before type Ia supernovae occur). These are indicated by ‘No’ in the third column. The other four occur alongside the CCSN explosion (or starting within seconds after the explosion).
@@
| AIC | [Co25]; [Pi26] | No | () Too long delay | () Too long delay |